# Splitting differential equations using Galois theory

### v1

@@ 1. Introduction
Differential algebra and model theory have a long history of fruitful interaction, starting with the work of Robinson [21], where differentially closed fields where introduced. Both disciplines inform each other: the theory of differentially closed fields is a rich source of examples in model theory, and model theoretic techniques can provide useful tools for the study of algebraic differential equations. A reason for this is $\omega$-stability of the theory $\mathrm{DCF}_{0}$ of differentially closed fields of characteristic zero, which brings all the machinery of finite rank stability theory into the picture.
@@
This article is concerned with one such tool, the semi-minimal analysis, and the development of Galois theoretic methods to control it, with an eye towards differential-algebraic applications. Recall that the semi-minimal analysis of a type $p\in S(A)$ is, roughly, a sequence of $A$-definable maps $p\xrightarrow{f_{1}}p_{1}\xrightarrow{f_{2}}\cdots\xrightarrow{f_{n}}p_{n}$ such that $p_{n}$, along with each fiber of an $f_{i}$, is semi-minimal, meaning internal to some rank one type. That this can be done in $\mathrm{DCF}_{0}$ gives a powerful way to decompose algebraic differential equations.
@@
To make this analysis useful, it is important to understand semi-minimal types. In the context of algebraic differential equations, this is done via Zilber’s dichotomy for $\mathrm{DCF}_{0}$: a semi-minimal type is either internal to a locally modular minimal type, and in particular has a rather rudimentary geometric structure, or is internal to the constants. We will be interested in this second possibility, of which we now give a geometrical account.
@@
A finite dimensional type $p$, over some algebraically closed differential field $k$, is always interdefinable with the generic type of an algebraic vector field $(V,s)$. Internality to the constants means that $(V,s)$ is, after possibly taking extensions to a differential field $k<L$, birationally equivalent to the trivial vector field $(\mathbb{A}^{\dim(V)},0)$. Equivalently, after a base change, there are $\dim(V)$ algebraically independent rational functions on $V$ that are constant for the induced derivation (but are not elements of $k$). These functions are called first integrals in the literature. Note that if we only ask for the existence of one such first integral, we obtain the notion of non-orthogonality to the constants. See [18] for more on this perspective.
@@
Even more concretely, the solution set of a differential equation (E) is internal to the constants if, roughly, there are finitely many solutions $a_{1},\cdots,a_{n}$ of (E) such that any other solution can be given as a rational function of the $a_{i}$ and finitely many constants. This is sometimes called a superposition principle in the literature (see [11] for example), and generalizes the fact that solution sets of linear differential equations are vectors spaces over the constant field.
@@
If equation (1) is internal to the constants, is equation (1) also internal to the constants?
@@
Note that because arbitrary field extensions are allowed, there is a priori no effective way to check if a given equation is internal to the constants. This is a general phenomenon in model-theoretic differential algebra: many useful concepts involve taking arbitrary field extensions, and controlling which extension one needs to look at has been an important theme. See for example recent work in [1] and [2] on the degree of non-minimality, which led to the development some effective tools for the study of algebraic differential equations.
Here, we will achieve some effectivity by considering a sufficient condition for internality, introduced in [10] and called splitting. Going back to the general case of a fibration $f:p\rightarrow f(p)$, we say that $f$ splits if there is in $A$-definable bijection $\iota$ between $p$ and the Morley product $f(p)\otimes r$, for some $\mathcal{C}$-internal type $r\in S(A)$, and the diagram
{tikzcd}
commutes. If we replace the bijection by a finite-to-finite correspondence, we obtain almost splitting. Note that there is no need to take extra parameters anymore. In the particular case of differential equations, it removes the need to take an arbitrary field extension. One of the pay-offs for understanding splitting is criteria for internality of algebraic differential equations, for example [10, Theorem B]. We will likewise produce such criteria here.
@@
Question 1 has mostly been asked for $m=1$ in the literature, but we remove that restriction here. One issue with picking $m>1$ (and working over possibly non-constant parameters) is that we loose a convenient criteria due to Rosenlicht [22] that tells us exactly when equation (1) is internal to the constants. To the authors’ knowledge, no such criteria is known without these assumptions.
To try to answer Question 1, we proceed in two steps, each of independent interest:
@@
Regarding step (b), to be more concrete, Equation 1 splits if, roughly, it can be transformed, using a differentially birational map, into an equation of the form:
$\begin{cases}\delta^{m}(y)=f(\delta^{m-1}(y),\cdots,\delta(y),y)\\ \delta(x)=\alpha x\end{cases}$
for some $\alpha\in k$. In the case $m=1$ and $k$ a field of constants, concrete algebraic characterizations of splitting were given in [5] and [10]. We extend them to arbitrary high order, and over a non-constant field:
@@
Let $F$ be an algebraically closed differential field, a rational function $f\in F(x_{0},\cdots,x_{m-1})$ and $q$ the generic type of $\delta^{m}(y)=f(y,\delta(y),\cdots,\delta^{m-1}(y))$. Let $p:=\log_{\delta}^{-1}(q)$ be the generic type of the pullback of $q$ under the logarithmic derivative.
Then $\log_{\delta}:p\rightarrow q$ is almost split if and only if there are a non-zero $h\in F(x_{0},\cdots,x_{m-1})$, some $e\in F$ and some integer $k\neq 0$ such that
@@
To understand splitting from a general, abstract perspective, we use model-theoretic Galois theory. Given a fixed definable set $\mathcal{C}$, any $\mathcal{C}$-internal type comes equipped with a definable group of transformations, called its binding group. In the case of algebraic differential equations, this corresponds to the Galois group associated to Kolchin’s strongly normal extensions [13], themselves a generalization of the Galois groups of Picard-Vessiot extensions. There is a Galois correspondence for these extensions, which was developed by Kolchin in [12] and [13].
To make use of binding groups, pick a fibration $f:p\rightarrow f(p)$, and assume that $p$ is $\mathcal{C}$-internal. This gives rise to a definable group morphism from the binding group of $p$ to the binding group of $f(p)$. Using the type-definable groupoids of [7], we connect splitting of the fibration $f$ to definable splitting of this map, obtaining necessary and sufficient conditions for splitting of $f$.
In the case of differential equations, binding groups over the constants are always definably isomorphic to the constant points of algebraic groups. This allows us to use classic classification results for algebraic groups to obtain some automatic splitting results, partially extending the results of [10]. We obtain:
Theorem B.
Let $F$ be an algebraically closed differential subfield of a monster model $\mathcal{U}\models\mathrm{DCF}_{0}$ and let $p\in S_{1}(F)$ be a $\mathcal{C}$-internal, weakly $\mathcal{C}$-orthogonal type with binding group definably isomorphic to the $\mathcal{C}$-points of a nilpotent algebraic group. Then $\log_{\delta}^{-1}(p)$ is almost $\mathcal{C}$-internal if and only if $\log_{\delta}:\log_{\delta}^{-1}(p)\rightarrow p$ almost splits.
Using non-split semi-abelian varieties, as well as a bit of algebraic group cohomology, we also answer a question in that same paper by showing the existence of a $\mathcal{C}$-internal type $p$ such that the map $\log_{\delta}:p\rightarrow\log_{\delta}(p)$ does not split, with the binding group of $\log_{\delta}(p)$ isomorphic to an elliptic curve:
@@
There exists an algebraically closed field $F<\mathcal{U}$, a type $q\in S_{1}(F)$ such that $\log_{\delta}:q\rightarrow\log_{\delta}(q)$ does not almost-split, the type $\log_{\delta}(q)$ is strongly minimal and has a binding group isomorphic to the constant points of an elliptic curve $E$.
@@
Section 3 consists of general model-theoretic results on splitting. We always consider a definable fibration $f:p\rightarrow f(p)$ with $\mathcal{C}$-internal fibers, for some fixed definable set $\mathcal{C}$. In Subsection 3.1, we define (almost) splitting of $f$ to be, roughly speaking, interdefinability (or interalgebraicity) between $p$ and $f(p)\otimes s$, for some $\mathcal{C}$-internal type $s$ over the same parameters. We connect it with uniform internality, as was done [5]. In Subsection 3.2, we introduce the main tool of our work: the connection between splitting of a fibration and splitting of a short exact sequence of definable groups. To do so, we make use of the type-definable groupoid associated to a fibration that was introduced in [7]. In particular, we use the notion of retractability of the groupoid. Roughly speaking, it means that we can uniformly and definably pick a family of maps between pairs of fibers of $f$, in a manner compatible with composition. We recall and expand on some results of that article, in particular connecting splitting of $f$ and retractability, but under the very restrictive the condition that $f(p)$ is fundamental. The main new result in that subsection is Lemma 3.16, which gives a sufficient condition to expand this when $f(p)$ is not fundamental: no subgroup of the binding group of $f(p)$ should map definably and with non-trivial image to the binding group of a fiber.
@@
In Section 4, we focus on the case of the logarithmic derivative. In subsection 4.1, we prove various results around splitting of the logarithmic derivative, generalizing a conjecture in Jin’s thesis [9, Conjecture 5.4]. This leads us to prove Theorem A (Lemma 4.8).
In Subsection 4.2, we consider any fibration $\log_{\delta}:p\rightarrow\log_{\delta}(p)$, where $p$ is internal, and the binding group of $\log_{\delta}(p)$ nilpotent. In that case, we obtain automatic splitting by Theorem B (Theorem 4.9). As an example of application, we prove that the logarithmic derivative pullback of the generic type of a Picard-Vessiot extension with nilpotent binding group is never almost internal to the constants.
Finally, Subsection 4.3 is dedicated to proving Theorem C (Corollary 4.19): we construct an example of a non-split fibration by the logarithmic derivative, answering a question of Jin and Moosa in [10, Bottom of page 5].
Acknowledgements. The authors are thankful to Rahim Moosa for multiple conversations on the subject of this article.
@@ 2.1. Internality and binding groups
In this article, we work, unless specified otherwise, in a sufficiently saturated model $\mathcal{U}$ of some totally transcendental theory $T$, eliminating imaginaries. Definable will always mean definable over some parameters, and we will always specify the parameters when relevant. When $p$ is a type, we denote $p(\mathcal{U})$ the set of realizations of $p$ in $\mathcal{U}$, and similarly for realizations of definable sets. We will often conflate definable sets and types with their set of realizations in $\mathcal{U}$. For this preliminary subsection, we fix an algebraically closed set of parameters $A$.
@@
Let $\mathcal{Q}$ be a family of partial types. A type $p\in S(A)$ is (almost) internal to $\mathcal{Q}$ is there are some $B\supset A$, some $a\models p$, some $c_{1},\cdots,c_{n}$ such that for all $i$ the type $\operatorname{tp}(c_{i}/B)$ extends a type in $\mathcal{Q}$, and:
@@
In practice, the minimal assumption for the theory to go through is that the family $\mathcal{Q}$ is $A$-invariant: any automorphism of $\mathcal{U}$ fixing $A$ fixes the set of realizations of $\mathcal{Q}$ pointwise. This happens, for example, if every partial type in $\mathcal{Q}$ is over $A$, which we assume for the rest of this subsection.
@@
Fundamental systems of solutions can be used to prove the following classical theorem (see [20, Theorem 7.4.8] for a proof):
@@
Let $\mathcal{Q}$ be a family of partial types over $A$, and $p\in S(A)$ a complete type, internal to $\mathcal{Q}$. Then $\mathrm{Aut}_{A}(p/\mathcal{Q})$ is isomorphic to an $A$-definable group, and its natural action on $p(\mathcal{U})$ is $A$-definable.
@@
  • orthogonal to $\mathcal{Q}$ if for any $B\supset A$, any realisation $a\models p|_{B}$ is independent, over $B$. from any tuple of realizations of $\mathcal{Q}$.
@@
Let $p\in S(A)$ be $\mathcal{C}$-internal. Then for any $B\supset A$, the binding group $\mathrm{Aut}_{B}(p|_{B}/\mathcal{C})$ is a definable normal subgroup of $\mathrm{Aut}_{A}(p/\mathcal{C})$.
@@
Pick a Morley sequence $\bar{a}=a_{1},\cdots,a_{n}$ of realizations of $p|_{B}$. It is also a Morley sequence in $p$, and we can pick $n$ large enough so that it is a fundamental system for both $p$ and $p|_{B}$. It is then easy to check that the map $\iota:\mathrm{Aut}_{B}(p|_{B}/\mathcal{C})\rightarrow\mathrm{Aut}_{A}(p/\mathcal{C})$ given by sending any $\sigma\in\mathrm{Aut}_{B}(p|_{B}/\mathcal{C})$ to the (unique) $\iota(\sigma)\in\mathrm{Aut}_{A}(p/\mathcal{C})$ such that $\iota(\sigma)(\bar{a})=\sigma(\bar{a})$ is definable and injective. It is straightforward to check that its image is normal. ∎
@@
Given two $\mathcal{C}$-internal types $p,q\in S(A)$, we can also form the binding groups $\mathrm{Aut}_{A}(p/\mathcal{C},q)$ and $\mathrm{Aut}_{A}(q/\mathcal{C},p)$. Then:
@@
We will make use of the following definable version of Goursat’s lemma (see [16, Chapter I, Exercise 5]), the proof of which is left to the reader:
@@
  • $p$ and $q$ are interdefinable if for any $a\models p$, there is $b\models q$ such that $\operatorname{dcl}(aA)=\operatorname{dcl}(bA)$. Note that by compactness, this implies that there is an $A$-definable bijection between $p(\mathcal{U})$ and $q(\mathcal{U})$.
@@
Note that if instead the types are interdefinable, the binding groups are isomorphic.

### v3

@@ 1. Introduction
Differential algebra and model theory have a long history of fruitful interaction, starting with the work of Robinson [25], where differentially closed fields were introduced. Both disciplines inform each other: the theory of differentially closed fields is a rich source of examples in model theory, and model theoretic techniques can provide useful tools for the study of algebraic differential equations. A reason for this is $\omega$-stability of the theory $\mathrm{DCF}_{0}$ of differentially closed fields of characteristic zero, which brings all the machinery of finite rank stability theory into the picture.
@@
This article is concerned with one such tool, the semi-minimal analysis, and the development of Galois theoretic methods to control it, with an eye towards differential-algebraic applications. Recall that a semi-minimal analysis of a type $p\in S(A)$ is a sequence of $A$-definable maps $p\xrightarrow{f_{1}}p_{1}\xrightarrow{f_{2}}\cdots\xrightarrow{f_{n}}p_{n}$ such that $p_{n}$, along with each fiber of an $f_{i}$, is semi-minimal, meaning internal to some rank one type. That this can be done in $\mathrm{DCF}_{0}$ gives a powerful way to decompose algebraic differential equations.
@@
To make this analysis useful, we need to understand semi-minimal types. For algebraic differential equations, this is done via Zilber’s dichotomy for $\mathrm{DCF}_{0}$: a semi-minimal type is either internal to a locally modular minimal type, and in particular has a rudimentary geometric structure, or is internal to the constants. We are interested in the second possibility, of which we give a geometrical account.
@@
A finite dimensional type $p$, over some algebraically closed differential field $k$, is always interdefinable with the generic type of a twisted algebraic vector field $(V,s)$, meaning that $s$ is a morphism over $k$ that is a section of the twisted tangent bundle (or prolongation) $\tau V$. In the autonomous case, i.e. when the derivative on $k$ is trivial, the twisted tangent bundle $\tau V$ is the tangent bundle $TV$. Internality to the constants means that $(V,s)$ is, after possibly taking extensions to a differential field $k<L$, birationally equivalent to the trivial vector field $(\mathbb{A}^{\dim(V)},0)$. We also have the close concept of almost internality: the twisted vector field in almost internal to the constants if, after a base change to some $k<K$, there are $\dim(V)$ rational functions on $V$, algebraically independent over $K$, that are constant for the induced derivation. These functions are called first integrals in the literature. Note that if we only ask for the existence of one such first integral, we obtain the notion of non-orthogonality to the constants. See [20] for more on this perspective.
@@
Even more concretely, the solution set of a differential equation (E) over some differential field $k$ is internal to the constants if there are finitely many solutions $a_{1},\cdots,a_{n}=\bar{a}$ of (E) and a rational function $h$ with coefficients in $k$ such that for any generic solution $a$ of (E), there are constants $c_{1},\cdots,c_{m}$ such that $a=h(\bar{a},\bar{a}^{\prime},\cdots,\bar{a}^{(l)})$ (where the derivative is taken coordinate-wise). This is sometimes called a superposition principle in the literature (see [13] for example), and generalizes the fact that solution sets of linear differential equations are vectors spaces over the constant field.
@@
If equation (1) is internal to the constants, is equation (1) also (almost) internal to the constants?
@@
Note that because arbitrary field extensions are allowed, there is a priori no effective way to check if a given equation is internal to the constants. This is a general phenomenon in model-theoretic differential algebra: many useful concepts involve taking arbitrary field extensions, and controlling which extension one needs to look at has been an important theme. See for example recent work in [3] and [4] on the degree of non-minimality, which led to the development of effective tools for the study of algebraic differential equations.
Here, we will achieve some effectivity by considering a sufficient condition for internality, introduced in [12] and which we will call splitting. Going back to the general case of a fibration $f:p\rightarrow f(p)$, we say that $f$ splits if there is an $A$-definable bijection $\iota$ between $p$ and the Morley product $f(p)\otimes r$, for some $\mathcal{C}$-internal type $r\in S(A)$, and the diagram
${p}$${f(p)\otimes r}$${f(p)}$$\scriptstyle{\iota}$$\scriptstyle{f}$
commutes. If we replace the bijection $\iota$ by a finite-to-finite correspondence, we obtain almost splitting. Note that there is no need to take extra parameters anymore. In the particular case of differential equations, it removes the need to take an arbitrary field extension. One of the pay-offs for understanding splitting is criteria for internality of algebraic differential equations, for example [12, Theorem B]. We will likewise produce such criteria here.
@@
Question 1 has been asked for $m=1$ in [12], but we remove that restriction here. One issue with picking $m>1$ (and working over possibly non-constant parameters) is that we loose a convenient criteria due to Rosenlicht [26] that tells us exactly when equation (1) is internal to the constants. To the authors’ knowledge, no such criteria is known without these assumptions  [^1] Since this article was first written, the authors [2] have obtained a generalization to $m>1$, over constant parameters..
To answer Question 1, we proceed in two steps, each of independent interest:
@@
For step (a), we use model-theoretic Galois theory, which makes sense in any $\omega$-stable theory. Given a fixed definable set $\mathcal{C}$, any $\mathcal{C}$-internal type comes equipped with a definable group of transformations, called its binding group. In the case of algebraic differential equations, this corresponds to the Galois group associated to Kolchin’s strongly normal extensions [14], themselves a generalization of the Galois groups of Picard-Vessiot extensions. As suggested by the anonymous referee, we now state what our methods yield in a general $\omega$-stable context, which is Theorem 3.18 below:
@@
Fix some $\omega$-stable theory $\mathrm{Th}$ eliminating imaginaries, some algebraically closed set of parameters $A$ and some $A$-definable pure algebraically closed field $\mathcal{C}$.
Let $p\in S(A)$ be a weakly $\mathcal{C}$-orthogonal type and $f:p\rightarrow q$ be an $A$-definable map. Assume that:
  • $q$ is $\mathcal{C}$-internal with binding group definably isomorphic to the $\mathcal{C}$-points of a linear nilpotent group and
  • for any $a\models p$, the fiber $\operatorname{tp}(a/f(a)A)$ is stationary and $\mathcal{C}$-internal with binding group isomorphic to the $\mathcal{C}$-points of a diagonalizable group.
Then $p$ is almost $\mathcal{C}$-internal if and only if the map $f$ almost splits.
In [12], Jin and Moosa prove a similar theorem in $\mathrm{DCF}_{0}$, for $q$ generic type of an equation of the form $\delta(x)=f(x)$ with $f\in F(X)$, internal to the constants and with binding group $G_{a}$, $G_{m}$ or $G_{a}\rtimes G_{m}$. Theorem A is therefore a generalization of their result, with the caveat that we now need the binding group to be linear nilpotent (thus excluding the $G_{a}\rtimes G_{m}$ case). The result should also hold for linear solvable binding groups, we reserve this question for future work.
To prove Theorem A, we make heavy use of binding groups. Consider a fibration $f:p\rightarrow f(p)$, and assume that $p$ is $\mathcal{C}$-internal. This gives rise to a definable group morphism from the binding group of $p$ to the binding group of $f(p)$. Using the type-definable groupoids of [9], we connect splitting of the fibration $f$ to definable splitting of the induced morphism between binding groups, thus obtaining necessary and sufficient conditions for splitting of $f$.
Because we are considering internality to a pure algebraically closed field $\mathcal{C}$, binding groups over the constants are always definably isomorphic to the constant points of algebraic groups. In particular, we obtain Theorem A by using the very constrained structure of nilpotent linear algebraic groups.
Regarding step (b), concrete characterizations of splitting were given in [12] and [7], if $k$ is a field of constants. Let $q$ be the generic type of the equation $\delta^{m}(y)=f(\delta^{m-1}(y),\cdots,\delta(y),y)$, splitting was characterized in the following cases:
  • in [12, Theorem B] for $m=1$ and $q$ internal to the constants,
  • in [7, Theorem 5.2], for arbitrary high $m$, but with $q$ orthogonal to the constants.
It follows from [7, Corollary 5.5] that the two criteria are in fact identical when $m=1$. This raises the possibility of a general characterization, free of assumptions on $m$ and $q$. We achieve that goal under a mild additional assumption:
Theorem B.
Let $F$ be an algebraically closed differential field, a rational function $f\in F(x_{0},\cdots,x_{m-1})$ and $q$ the generic type of $\delta^{m}(y)=f(y,\delta(y),\cdots,\delta^{m-1}(y))$. Let $p:=\log_{\delta}^{-1}(q)$, and suppose that $p(\mathcal{U})\not\subset\operatorname{acl}(q(\mathcal{U}),\mathcal{C},F)$. Then $\log_{\delta}:p\rightarrow q$ is almost split if and only if there are a non-zero $h\in F(x_{0},\cdots,x_{m-1})$, some $e\in F$ and some integer $k\neq 0$ such that:
$(kx_{0}-e)h=\sum\limits_{i=0}^{m-2}\frac{\partial h}{\partial x_{i}}x_{i+1}+\frac{\partial h}{\partial x_{m-1}}f+\delta^{F}(h)$
@@
This theorem can be considered a simultaneous generalization of the criteria given by [7, Corollary 5.5] and [12, Theorem B], under the extra assumption that $p(\mathcal{U})\not\subset\operatorname{acl}(\log_{\delta}(p)(\mathcal{U}),\mathcal{C},F)$. Let us say a few words about this assumption. The proof of Theorem B goes through the auxiliary notion of product-splitting: we first show that almost splitting is equivalent to product splitting, and then show that product splitting is equivalent to the second, algebraic property of Theorem B. To deduce product-splitting from almost splitting, we need the binding group of the type $r$ witnessing almost splitting to be infinite. The extra assumption is the weakest we could find that implies this. Note that we can prove that the binding group of $r$ is infinite under other extra assumptions, such as when $q$ is orthogonal to the constants (Corollary 4.7), or $F$ is a field of constants (Corollary 4.6).
Together, Theorem A and Theorem B allow us to prove that some pullbacks under the logarithmic derivative are not internal: we first show they must be split using Theorem A, and then use the concrete algebraic characterization of splitting given by Theorem B to prove that splitting is impossible. As an example of application, we prove that if $a\models q$ is a generic solution of a linear differential equation, then $\log_{\delta}^{-1}(q)$ is never $\mathcal{C}$-internal, provided that the binding group of $q$ is nilpotent and acts transitively on $q$.
Finally, our methods provide examples of internal types $p$ with a non-split fibration $f:p\rightarrow f(p)$. More precisely, using non-split semi-abelian varieties, as well as a bit of algebraic group cohomology, we also answer a question of Jin and Moosa in [12] by showing the existence of a $\mathcal{C}$-internal type $p$ such that the map $\log_{\delta}:p\rightarrow\log_{\delta}(p)$ does not split, with the binding group of $\log_{\delta}(p)$ isomorphic to an elliptic curve:
@@
There exist an algebraically closed field $F<\mathcal{U}$ and a $\mathcal{C}$-internal type $p\in S_{1}(F)$ such that $\log_{\delta}:p\rightarrow\log_{\delta}(p)$ does not almost-split, and the type $\log_{\delta}(p)$ is strongly minimal and has a binding group isomorphic to the constant points of an elliptic curve $E$.
In particular, Theorem C shows that in Theorem A, it is necessary to assume that the binding group of the base $q$ is linear.
@@
Section 3 consists of general model-theoretic results on splitting. We always consider a definable fibration $f:p\rightarrow f(p)$ with $\mathcal{C}$-internal fibers, for some fixed definable set $\mathcal{C}$. In Subsection 3.1, we define splitting of $f$ and connect it to uniform internality, and prove a basic result we will need. In Subsection 3.2, we introduce the main tool of our work: the connection between splitting of a fibration and splitting of a short exact sequence of definable groups. To do so, we make use of the type-definable groupoid associated to a fibration that was introduced in [9]. In particular, we use the notion of retractability of the groupoid introduced in [5], also reminiscent of descent arguments, see for example [23, Proposition 4.2]. If $\mathcal{G}$ is a groupoid, it is retractable if there is a definable map $r:\mathrm{Ob}(\mathcal{G})^{2}\rightarrow\mathrm{Mor}(\mathcal{G})$ such that for all $a,b,c\in\mathrm{Ob}(\mathcal{G})$, we have $r(a,b)\in\mathrm{Mor}(a,b)$ and $r(b,c)\circ r(a,b)=r(a,c)$. We recall and expand on some results of that article, in particular connecting splitting of $f$ and retractability, but under the very restrictive the condition that $f(p)$ is fundamental. The main new result on groupoids is Lemma 3.14, which gives a sufficient condition to expand this when $f(p)$ is not fundamental: no subgroup of the binding group of $f(p)$ should map definably and with non-trivial image to the binding group of a fiber. We end the section by stating and proving Theorem A.
@@
In Section 4, we focus on the case of the logarithmic derivative. In subsection 4.1, we prove various results around splitting of the logarithmic derivative. We connect it with another notion of splitting, which we call product-splitting, introduced by Jin in his thesis [11] and further studied by Jin and Moosa in [12]. In [11, Conjecture 5.4], Jin conjectures that almost splitting and product splitting are equivalent in the specific case where the image of the fibration is internal. We prove that splitting and product-splitting are equivalent under some various additional assumptions, making progress on Jin’s conjecture. We then prove an algebraic characterization of product-splitting, which leads us to prove Theorem B (Corollary 4.9).
In Subsection 4.2, we consider any fibration $\log_{\delta}:p\rightarrow\log_{\delta}(p)$, where $p$ is almost internal to the constants, and the binding group of $\log_{\delta}(p)$ nilpotent. In that case, we obtain automatic splitting by Theorem A (Theorem 4.10). As an example of application, we prove that the logarithmic derivative pullback of the generic type of a linear differential equation with nilpotent binding group is never almost internal to the constants.
Finally, Subsection 4.3 is dedicated to proving Theorem C (Corollary 4.20): we construct an example of a non-split fibration by the logarithmic derivative, answering a question of Jin and Moosa in [12, bottom of page 5].
Acknowledgements. The authors are thankful to Rahim Moosa for multiple conversations on the subject of this article. We are also grateful to the anonymous referee, whose careful reading of our work lead to numerous improvements, in particular showing the full equivalence between almost internality and almost splitting in Theorem A.
@@ 2.1. Internality and binding groups
In this article, we work, unless specified otherwise, in a sufficiently saturated model $\mathcal{U}$ of some $\omega$-stable theory $\mathrm{Th}$, eliminating imaginaries. We will make use, without mentioning it, of the fact that in this context, type-definable groups are always definable (see [24, Corollary 5.19]). Definable will always mean definable over some parameters, and we will always specify the parameters when relevant. When $p$ is a type, we denote $p(\mathcal{U})$ the set of realizations of $p$ in $\mathcal{U}$, and similarly for realizations of definable sets. We will often conflate definable sets and types with their set of realizations in $\mathcal{U}$. For example, we may say that that a definable group $G$ acts on a type $p$ to mean that it acts on $p(\mathcal{U})$. We fix a set of parameters $A$.
@@
Let $\mathcal{Q}$ be a family of partial types. A stationary type $p\in S(A)$ is (almost) internal to $\mathcal{Q}$ if there are some $B\supset A$, some $a\models p$, some $c_{1},\cdots,c_{n}$ such that for all $i$ the type $\operatorname{tp}(c_{i}/B)$ extends a type in $\mathcal{Q}$, and:
@@
In practice, the minimal assumption for the theory to go through is that the family $\mathcal{Q}$ is $A$-invariant: any automorphism of $\mathcal{U}$ fixing $A$ fixes the set of realizations of $\mathcal{Q}$ setwise. This happens, for example, if every partial type in $\mathcal{Q}$ is over $A$, which we assume for the rest of this subsection.
@@
Fundamental systems of solutions can be used to prove the following classical theorem (see [22, Theorem 7.4.8] for a proof):
@@
Let $\mathcal{Q}$ be a family of partial types over $A$, and $p\in S(A)$ a complete type, internal to $\mathcal{Q}$. Then $\mathrm{Aut}_{A}(p/\mathcal{Q})$ is isomorphic to an $A$-definable group, and its natural action on $p(\mathcal{U})$ is relatively $A$-definable.
@@
  • orthogonal to $\mathcal{Q}$ if for any $B\supset A$, any realisation $a\models p|_{B}$ is independent, over $B$, from any tuple of realizations of $\mathcal{Q}$.
@@
Let $p\in S(A)$ be $\mathcal{C}$-internal. Then for any $B\supset A$, the binding group $\mathrm{Aut}_{B}(p|_{B}/\mathcal{C})$ is a definable subgroup of $\mathrm{Aut}_{A}(p/\mathcal{C})$.
@@
Pick a Morley sequence $\bar{a}=a_{1},\cdots,a_{n}$ of realizations of $p|_{B}$. It is also a Morley sequence in $p$, and we can pick $n$ large enough so that it is a fundamental system for both $p$ and $p|_{B}$. It is then easy to check that the map $\iota:\mathrm{Aut}_{B}(p|_{B}/\mathcal{C})\rightarrow\mathrm{Aut}_{A}(p/\mathcal{C})$ given by sending any $\sigma\in\mathrm{Aut}_{B}(p|_{B}/\mathcal{C})$ to the (unique) $\iota(\sigma)\in\mathrm{Aut}_{A}(p/\mathcal{C})$ such that $\iota(\sigma)(\bar{a})=\sigma(\bar{a})$ is definable and injective. ∎
@@
Given two types $p,q\in S(A)$ such that $p$ is $\mathcal{C}$-internal, we can also form the binding group $\mathrm{Aut}_{A}(p/\mathcal{C},q)$. Then:
@@
We will make use of the following definable version of Goursat’s lemma (see [18, Chapter I, Exercise 5]), the proof of which is left to the reader:
@@
  • $p$ and $q$ are interdefinable if for any $a\models p$, there is $b\models q$ such that $\operatorname{dcl}(aA)=\operatorname{dcl}(bA)$. Note that by compactness, this is equivalent to the existence of an $A$-definable bijection between $p(\mathcal{U})$ and $q(\mathcal{U})$.
@@
If instead the types are interdefinable, the binding groups are isomorphic.